OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} ( 1+x+f(x)/x) ) ^{1/x} = e^3\] then find \[\lim_{x \rightarrow 0} \frac{ f(x) }{ x^2}\]

OpenStudy (anonymous):

@satellite73 @Mertsj @zepdrix @eashmore @bahrom7893

OpenStudy (zarkon):

looks like it would be 2

OpenStudy (anonymous):

Yeah ur answer is correct can u Xplain?

OpenStudy (anonymous):

@amistre64 @phi @AravindG

OpenStudy (aravindg):

Do you know the general expression lim x-->0 (1+ax)^{1/x} ?

OpenStudy (anonymous):

taking log ryt?

OpenStudy (aravindg):

Nope. Its ready made formula. It equals e^a

OpenStudy (aravindg):

I hope you get the idea from that.

OpenStudy (zarkon):

Let \(g(x)=\dfrac{f(x)}{x}\) then you want \[\lim_{x\to0}\frac{g(x)}{x}\] take the log of the original and you have \[\lim_{x\to 0}\frac{\ln(1+x+g(x))}{x}=3\] use L'hospitals rule \[3=\lim_{x\to 0}\frac{\ln(1+x+g(x))}{x}=\lim_{x\to 0}\frac{1+g'(x)}{1+x+g(x)}\] clearly \(1+x+g(x)\to 1\) otherwise the original limit would not be exp(3) so we have \(3=\displaystyle\lim_{x\to 0}(1+g'(x))\) so \[\lim_{x\to 0}g'(x)=2\] then \[\lim_{x\to0}\frac{f(x)}{x^2}=\lim_{x\to0}\frac{g(x)}{x}=\lim_{x\to 0}\frac{g'(x)}{1}=2\]

OpenStudy (phi):

Here is another way to do this. According to https://en.wikipedia.org/wiki/Exponential_function#Formal_definition \[e^x= \lim_{n \rightarrow \infty} \left( 1+\frac{ x }{ n } \right)^n\] replace x with 3 and do a change of variables n =1/x to get \[ e^3= \lim_{x \rightarrow 0} \left( 1+3x \right)^\frac{1}{x}\] pattern match this with your problem \[ \lim_{x \rightarrow 0} \left( 1+x+\frac{f(x)}{x}\right ) ^{1/x} = e^3 \] and we see that \[ \lim_{x \rightarrow 0}\left( x+\frac{f(x)}{x}=3x\right )\] divide by x : \[ \lim_{x \rightarrow 0}\left( 1+\frac{f(x)}{x^2}=3\right ) \\ \lim_{x \rightarrow 0}\left( \frac{f(x)}{x^2}=2\right ) \\ \lim_{x \rightarrow 0}\frac{f(x)}{x^2}=2 \]